# NewtonBench: Benchmarking Generalizable Scientific Law Discovery in LLM Agents

### v1

@@ B.2 Results of Non-Performance Metrics
GPT-4.1-mini 6.88 6.97 7.53 6.38 6.50 6.74 6.40 6.65 6.87 6.77
GPT-4.1 8.01 8.52 8.72 7.44 7.62 7.45 7.09 7.26 7.57 7.74
o4-mini 2.83 3.18 3.80 2.94 3.40 3.83 2.99 3.19 4.10 3.36
GPT-5-mini 2.31 2.51 3.20 2.18 2.41 2.97 2.76 2.76 3.13 2.69
GPT-5 2.40 2.38 2.80 2.56 2.65 3.39 3.17 3.26 4.02 2.96
DeepSeek-V3 4.62 4.88 5.08 4.18 4.35 4.51 4.24 4.49 4.47 4.53
DeepSeek-R1 3.03 3.35 4.08 3.40 3.47 4.15 3.72 3.60 4.23 3.67
QwQ-32b 3.03 3.60 4.64 3.16 3.27 3.79 2.99 2.85 3.40 3.41
Qwen-3-235b 2.92 3.22 3.49 3.00 2.92 3.25 3.53 3.53 3.76 3.29
Gemini-2.5-flash 4.49 4.95 5.81 4.43 5.10 5.76 5.35 5.53 6.33 5.31
Gemini-2.5-pro 3.03 3.06 3.11 2.95 2.88 3.15 3.32 3.10 3.64 3.14
@@
GPT-4.1-mini 8.11 8.81 8.95 8.21 8.75 8.96 8.58 8.78 8.86 8.67
GPT-4.1 7.85 10.03 10.92 9.44 10.23 10.66 10.16 10.71 10.76 10.08
o4-mini 4.22 5.20 6.90 4.53 5.67 6.43 4.85 5.59 6.22 5.51
GPT-5-mini 4.12 5.18 5.99 4.51 5.83 6.55 5.31 5.90 5.97 5.48
GPT-5 3.60 4.15 5.38 4.27 5.02 6.45 5.20 5.76 6.97 5.20
DeepSeek-V3 7.85 9.20 8.41 7.76 8.29 7.30 8.10 8.00 7.44 8.04
DeepSeek-R1 5.12 5.46 5.92 5.28 5.48 5.97 5.69 5.78 6.20 5.66
QwQ-32b 3.95 6.00 5.94 4.24 5.94 5.01 4.72 4.88 4.97 5.07
Qwen-3-235b 4.60 4.97 5.42 4.98 5.28 5.35 5.22 5.49 5.74 5.23
Gemini-2.5-flash 7.24 8.60 9.73 7.69 8.64 9.90 8.27 8.83 9.39 8.70
Gemini-2.5-pro 5.51 5.44 6.46 5.45 5.77 6.68 6.03 6.26 7.08 6.08
@@
GPT-4.1-mini 44.78 43.60 47.65 30.98 32.23 33.96 33.25 34.81 36.26 37.50
GPT-4.1 105.86 111.17 113.44 69.00 72.51 73.47 63.21 63.23 66.40 82.03
o4-mini 17.23 19.35 26.05 13.48 17.10 21.95 12.67 14.85 21.57 18.25
GPT-5-mini 19.80 22.74 31.81 14.47 18.15 24.83 21.23 21.00 26.80 22.31
GPT-5 24.12 24.07 30.21 22.98 23.42 35.81 33.51 34.87 44.92 30.43
DeepSeek-V3 22.05 24.03 24.51 14.74 15.88 16.98 14.94 15.87 15.83 18.31
DeepSeek-R1 22.37 24.49 28.32 17.50 18.58 22.37 17.49 17.58 18.14 20.76
QwQ-32b 13.29 13.29 13.12 8.44 7.22 10.08 6.93 7.22 8.16 9.75
Qwen-3-235b 9.12 9.83 10.76 6.85 6.51 7.72 7.59 7.34 8.13 8.21
Gemini-2.5-flash 21.99 25.49 30.79 20.23 23.77 27.06 19.98 22.59 26.25 24.24
Gemini-2.5-pro 16.31 16.48 17.85 13.40 13.28 15.08 15.74 14.22 17.47 15.54
@@
GPT-4.1-mini 21.11 22.10 24.55 14.36 15.70 16.40 15.51 16.13 18.78 18.29
GPT-4.1 30.22 38.51 44.89 27.82 28.11 30.49 25.47 25.01 26.80 30.81
o4-mini 13.74 16.01 22.26 10.88 13.83 17.35 10.15 13.39 15.13 14.75
GPT-5-mini 16.11 21.29 27.49 15.22 20.58 25.22 17.61 19.74 24.35 20.84
GPT-5 20.40 21.72 27.24 19.54 22.40 28.45 25.98 26.12 33.03 24.99
DeepSeek-V3 15.06 18.24 18.11 12.71 13.40 12.42 12.07 11.92 11.18 13.90
DeepSeek-R1 12.63 15.47 16.20 11.06 13.42 13.30 10.98 11.26 10.35 12.74
QwQ-32b 9.59 10.67 10.99 6.69 6.33 6.80 6.28 5.90 5.85 7.68
Qwen-3-235b 8.19 8.50 8.73 6.43 6.93 6.59 5.99 6.08 6.98 7.16
Gemini-2.5-flash 15.51 15.49 17.35 12.72 13.16 14.67 13.10 13.84 15.69 14.61
Gemini-2.5-pro 12.56 11.47 12.41 9.42 9.44 10.38 10.22 9.02 11.68 10.73
@@
GPT-4.1-mini 1284 1591 1521 1610 1944 2612 1856 1711 2096 1803
GPT-4.1 2605 2790 2810 2589 2832 2898 2911 3324 3324 2898
o4-mini 4624 9675 19944 7254 12149 17196 9115 11619 18918 12277
GPT-5-mini 3457 7574 15458 5252 10009 14396 8021 10589 13924 9853
GPT-5 7101 8372 19424 11214 16862 31388 17054 22722 37455 19066
DeepSeek-V3 1516 1582 1646 1543 1693 1665 1525 1607 1541 1591
DeepSeek-R1 9724 13538 18337 14647 17972 19666 16751 18260 19231 16458
QwQ-32b 9450 13422 17461 11947 13945 16716 12838 14326 15814 13991
Qwen-3-235b 8104 12615 16282 10729 12944 14313 12603 13517 14616 12858
Gemini-2.5-flash 8537 19327 40681 18696 32575 42013 38114 32682 49863 31388
Gemini-2.5-pro 9238 13721 23950 13961 18215 28272 22203 24319 33078 20773
@@
GPT-4.1-mini 1919 2312 2465 2877 2989 3371 3635 3579 3523 2963
GPT-4.1 2989 4798 5836 4259 4987 5181 4774 4914 5419 4795
o4-mini 4423 9648 18410 6214 11193 15168 8453 11280 14227 11002
GPT-5-mini 3702 8748 12927 6790 11643 13660 10213 11399 12514 10177
GPT-5 4821 9168 20185 10678 17893 30240 17431 23295 35488 18800
DeepSeek-V3 2148 2593 2536 2454 2692 2345 2684 2589 2449 2499
DeepSeek-R1 11281 14616 18596 15372 17651 19815 18014 17719 18677 16860
QwQ-32b 7724 13689 17667 10872 13985 16915 13638 14177 15890 13840
Qwen-3-235b 6881 11672 15832 10516 12612 14301 12506 13960 14114 12488
Gemini-2.5-flash 5388 18256 24028 10225 18092 21096 19407 27087 25411 18777
Gemini-2.5-pro 9301 12932 25447 13820 19348 30906 22930 29170 36998 22317
@@
GPT-4.1-mini 198 272 216 309 415 414 319 270 366 309
GPT-4.1 335 341 329 367 408 421 463 501 488 406
o4-mini 1633 2771 4694 2257 3318 4107 2918 3420 4321 3271
GPT-5-mini 1518 2825 4490 2365 3890 4490 2925 3780 4416 3411
GPT-5 2599 3367 6168 4005 5937 8657 5004 6344 8561 5627
DeepSeek-V3 379 366 356 415 426 400 412 394 399 394
DeepSeek-R1 3351 4294 4837 4612 5771 5179 4991 5523 5190 4861
QwQ-32b 3711 4703 4999 4756 5575 5510 5754 6364 6151 5280
Qwen-3-235b 3148 4270 5284 4391 5081 5103 4253 4628 4795 4550
Gemini-2.5-flash 1941 3889 6904 3982 5879 7591 6090 6038 7563 5542
Gemini-2.5-pro 3124 4546 7757 4806 6409 9201 7026 8359 9531 6751
@@
GPT-4.1-mini 233 260 272 355 343 378 424 403 398 341
GPT-4.1 360 454 517 454 493 490 474 472 508 469
o4-mini 983 1684 2495 1264 1794 2191 1609 1877 2101 1778
GPT-5-mini 870 1538 2046 1426 1901 2017 1795 1923 2057 1730
GPT-5 1208 1972 3264 2279 3266 4363 3130 3803 4857 3127
DeepSeek-V3 268 280 303 317 334 325 342 333 341 316
DeepSeek-R1 2281 2793 3340 3092 3405 3568 3279 3211 3123 3121
QwQ-32b 2043 2640 3436 2754 3023 3698 3280 3309 3684 3096
Qwen-3-235b 1568 2406 3039 2246 2516 2820 2621 2814 2680 2524
Gemini-2.5-flash 709 1656 2413 1297 2193 2221 2361 3162 2828 2093
Gemini-2.5-pro 1672 2344 3797 2493 3251 4671 3668 4703 5164 3529
@@ E.1 Memorization in LSR-Transform
With the enhanced reasoning abilities of frontier LLMs, the effectiveness of equation transformation in preventing LLM memorization may vary. In Table 11, we present the performance of eight LLMs on LSR-Transform (Shojaee et al., 2025b) without observing any data. We observe that the latest LLMs demonstrate significant improvements over previous, weaker models, indicating their ability to deduce the transformed law from the original solely based on memorized knowledge (with an example from o4-mini below). On the other hand, as NewtonBench curates counterfactual physical laws via metaphysical shift, making recall of canonical laws insufficient. Therefore, all LLMs achieve zero scores in the DataBlind setting.
@@
That is,
m = 4 E_n / [x² (omega² + omega_0²)].
@@ E.4 Exploration and Exploitation Tokens
E.5 Use of Large Language Models

### v2

@@ B.2 Results of Non-Performance Metrics
GPT-4.1-mini 7.00 6.89 6.90 6.49 6.64 6.56 6.54 6.79 6.85 6.74
GPT-4.1 7.89 8.59 8.63 7.22 7.42 7.35 6.99 7.12 7.19 7.60
o4-mini 2.74 2.97 3.53 2.85 3.21 3.82 3.23 3.58 4.25 3.35
GPT-5-mini 2.19 2.36 2.88 2.16 2.40 2.76 2.78 2.90 3.03 2.60
GPT-5 2.41 2.38 2.60 2.60 2.65 3.19 3.13 3.40 3.89 2.92
DeepSeek-V3 4.49 4.97 5.15 4.21 4.42 4.62 4.28 4.49 4.33 4.55
DeepSeek-R1 2.92 3.24 3.85 3.15 3.35 3.84 3.73 3.56 4.04 3.52
QwQ-32b 2.80 2.97 3.64 2.75 2.72 3.17 2.74 2.46 2.95 2.91
Qwen-3-235b 3.22 3.38 3.62 3.12 3.26 3.40 4.01 3.85 4.04 3.54
Gemini-2.5-flash 3.99 4.67 5.43 4.43 4.62 5.51 5.47 5.65 6.33 5.12
Gemini-2.5-pro 2.99 3.06 3.08 3.00 2.92 3.07 3.36 3.16 3.47 3.12
@@
GPT-4.1-mini 8.10 8.56 8.77 8.52 8.62 8.90 8.61 8.94 8.94 8.66
GPT-4.1 8.02 8.82 9.65 9.08 9.60 9.96 9.51 10.25 10.25 9.46
o4-mini 4.28 4.99 6.40 4.60 5.30 6.13 5.05 5.92 6.66 5.48
GPT-5-mini 4.01 4.83 5.51 4.61 5.25 5.91 5.17 5.51 5.94 5.19
GPT-5 3.74 4.28 5.33 4.35 5.08 6.35 5.36 5.90 7.13 5.28
DeepSeek-V3 6.93 7.77 7.53 6.62 7.04 6.86 6.90 6.92 7.03 7.07
DeepSeek-R1 5.25 5.50 5.81 5.40 5.63 6.08 5.81 5.94 6.37 5.75
QwQ-32b 3.72 4.62 4.47 3.93 4.22 4.01 4.06 3.88 4.21 4.12
Qwen-3-235b 4.60 5.16 5.41 5.01 5.51 5.37 5.39 5.51 5.84 5.31
Gemini-2.5-flash 7.08 8.35 9.03 7.90 8.56 9.12 8.24 8.72 9.03 8.45
Gemini-2.5-pro 5.45 5.68 6.26 5.45 5.98 6.77 6.17 6.40 7.40 6.17
@@
GPT-4.1-mini 44.78 42.48 41.40 32.37 33.89 32.07 35.49 36.64 37.27 37.38
GPT-4.1 104.70 109.40 110.08 70.47 71.97 74.01 60.66 61.22 62.79 80.59
o4-mini 16.78 18.45 23.76 13.21 15.66 21.15 14.97 17.13 21.22 18.04
GPT-5-mini 19.05 20.60 27.07 14.42 17.96 21.96 21.38 22.61 24.52 21.06
GPT-5 24.29 23.99 26.74 23.18 23.35 31.82 32.72 35.74 41.39 29.25
DeepSeek-V3 21.40 23.72 24.53 14.99 15.99 17.41 14.63 15.81 15.28 18.20
DeepSeek-R1 21.63 23.70 27.12 17.26 18.81 21.44 17.81 16.96 18.00 20.30
QwQ-32b 12.94 13.02 12.28 7.76 7.60 9.74 6.77 5.74 7.48 9.26
Qwen-3-235b 9.01 9.95 10.60 6.90 6.83 6.97 7.81 7.47 8.08 8.18
Gemini-2.5-flash 22.44 25.52 29.17 20.19 21.35 23.47 21.08 22.65 24.81 23.41
Gemini-2.5-pro 15.90 16.01 16.94 13.52 13.51 14.17 15.98 15.46 15.84 15.26
@@
GPT-4.1-mini 21.03 20.83 24.00 15.27 14.62 15.17 16.45 16.08 18.12 17.95
GPT-4.1 31.55 34.74 39.69 30.48 29.69 31.41 25.84 27.52 27.19 30.90
o4-mini 13.61 14.85 18.60 10.92 11.96 14.98 11.08 12.97 15.24 13.80
GPT-5-mini 16.39 20.12 24.03 15.47 18.44 21.84 17.51 19.81 23.70 19.70
GPT-5 20.59 21.20 24.21 20.06 21.94 26.08 26.42 27.37 33.02 24.54
DeepSeek-V3 14.11 15.74 16.28 10.61 11.52 11.90 10.41 10.98 11.15 12.52
DeepSeek-R1 12.49 15.01 16.60 11.10 12.00 12.94 10.92 11.49 10.35 12.54
QwQ-32b 9.03 10.27 9.81 6.35 6.01 6.22 5.38 5.27 5.39 7.08
Qwen-3-235b 7.69 8.10 8.37 6.06 6.88 6.25 5.72 5.33 6.03 6.71
Gemini-2.5-flash 14.95 15.83 16.26 13.74 13.78 13.16 12.06 14.49 15.00 14.36
Gemini-2.5-pro 12.33 11.42 11.47 9.18 9.40 9.37 10.65 9.92 12.00 10.64
@@
GPT-4.1-mini 1213 1522 2333 1492 2126 3689 1726 1691 2002 1977
GPT-4.1 2510 2588 2660 2623 2808 2809 2897 3115 3371 2820
o4-mini 4487 7629 15426 6695 9924 16613 10199 13326 19836 11571
GPT-5-mini 3508 6348 12236 5289 9188 13324 9301 11595 13915 9412
GPT-5 7397 9055 18334 12138 17479 29562 17730 24577 36035 19145
DeepSeek-V3 1562 1530 1738 1509 1672 1688 1487 1545 1554 1587
DeepSeek-R1 9733 13672 17530 14485 18100 19442 17203 18293 19687 16461
QwQ-32b 10335 13658 16725 12293 14827 18079 13841 14239 16415 14490
Qwen-3-235b 8560 12622 16366 12045 13190 16106 14653 14875 16736 13906
Gemini-2.5-flash 8702 21702 37980 18832 28682 38817 43646 33627 57040 32114
Gemini-2.5-pro 8823 15748 24543 14105 19246 26718 24380 26864 33422 21539
@@
GPT-4.1-mini 1906 2196 2322 2872 2920 3140 3565 3613 3701 2915
GPT-4.1 3030 3798 4814 4465 4728 5054 4518 5127 5267 4534
o4-mini 4310 7996 14774 6372 8983 12678 8948 12144 16121 10258
GPT-5-mini 3552 6964 9754 7029 9401 11778 9983 12034 13152 9294
GPT-5 5086 9034 17016 10747 16368 27063 19395 25450 35032 18355
DeepSeek-V3 1957 2143 2256 2107 2274 2269 2355 2333 2342 2226
DeepSeek-R1 10847 14842 17489 15566 18222 19245 17934 18490 19619 16917
QwQ-32b 7999 13668 17719 12156 15048 16889 15698 15246 17171 14621
Qwen-3-235b 6887 12136 16007 11303 13835 15822 13589 14268 15871 13302
Gemini-2.5-flash 5532 14134 15635 11682 13871 18230 19702 32988 27757 17726
Gemini-2.5-pro 9080 13219 20955 13803 18559 29463 24979 29074 38242 21930
@@
GPT-4.1-mini 181 259 442 276 431 699 290 263 353 355
GPT-4.1 337 315 319 403 418 424 470 492 532 412
o4-mini 1628 2537 3985 2128 2906 4010 3082 3684 4507 3163
GPT-5-mini 1597 2671 4035 2399 3717 4682 3299 4024 4670 3455
GPT-5 2718 3609 6277 4329 6203 8587 5482 6934 8996 5904
DeepSeek-V3 412 340 350 414 429 406 401 368 409 392
DeepSeek-R1 3476 4453 4955 4986 5992 5702 5068 5753 5591 5108
QwQ-32b 4226 5399 6073 5485 6609 6664 6751 7464 7410 6231
Qwen-3-235b 3182 4204 5210 4617 4874 5547 4352 4774 5126 4654
Gemini-2.5-flash 2110 4112 6650 3832 5774 7345 7064 6035 8660 5731
Gemini-2.5-pro 3013 5143 7941 4760 6664 9014 7386 8759 10217 6988
@@
GPT-4.1-mini 231 254 261 342 342 356 413 403 416 335
GPT-4.1 354 406 480 493 500 516 475 510 522 473
o4-mini 953 1503 2126 1279 1558 1936 1645 1941 2277 1691
GPT-5-mini 857 1325 1660 1458 1707 1936 1816 2122 2190 1674
GPT-5 1242 1907 2755 2277 3003 3979 3465 4180 4942 3083
DeepSeek-V3 273 274 299 315 328 335 343 340 349 317
DeepSeek-R1 2122 2847 3221 3094 3495 3435 3229 3312 3221 3108
QwQ-32b 2245 3229 4209 3265 3785 4402 4221 4340 4567 3807
Qwen-3-235b 1580 2384 3057 2435 2741 3120 2777 2818 2984 2655
Gemini-2.5-flash 744 1613 1767 1504 1651 2131 2405 3873 3198 2099
Gemini-2.5-pro 1643 2219 3272 2514 2983 4336 3852 4547 5227 3399
@@ Appendix D Case Studies
NewtonBench introduces an active exploration paradigm to enable LLM agents to interact with virtual environments and discover novel physical laws. Figure 20 intuitively illustrates the paradigm difference between the previous passive observation paradigm and our active exploration paradigm.
Figure 20: An illustration of the active exploration paradigm presented in NewtonBench, in comparison to passive observation paradigm in previous symbolic regression benchmarks.
@@ E.1 Memorization in LSR-Transform
With the enhanced reasoning abilities of frontier LLMs, the effectiveness of equation transformation in preventing LLM memorization may vary. In Table 17, we present the performance of eight LLMs on LSR-Transform (shojaee2025llmsrbenchnewbenchmarkscientific) without observing any data. We observe that the latest LLMs demonstrate significant improvements over previous, weaker models, indicating their ability to deduce the transformed law from the original solely based on memorized knowledge (with an example from o4-mini below).
On the other hand, as NewtonBench curates counterfactual physical laws via counterfactual shift, making recall of canonical laws insufficient. To further verify that performance on NewtonBench does not rely on memorization, we conduct a DataBlind experiment in the same setting, where models are required to generate hypotheses purely from contextual information without access to any observed data; the results are reported in Table 18. We find that, even when given an explicit hint indicating that the target law is shifted, all models still achieve zero accuracy (with an example from GPT-4.1 below), which provides additional evidence that NewtonBench is effectively memorization-free.
@@
Table 18: DataBlind performance on NewtonBench
Models Accuracy w/o hint Accuracy with hint
GPT-4o-mini 0 0
Nemotron-ultra-253b 0 0
GPT-4.1 0 0
@@
That is, m = 4 E_n / [x² (omega² + omega_0²)].
NewtonBench Example Question
universe. Please note that the laws of physics in this universe are different from those
in our own. You will be given a short textual description of the target law (inputs and
expected physical relationship). You have no example numeric data. Based on the
description alone, provide your best guess of the law.
You need to discover an altered Newton’s Law of Universal Gravitation. Inputs:
- mass1: mass of the first object. It should be a positive real number.
- mass2: mass of the second object. It should be a positive real number.
- distance: distance between the two objects. It should be a positive real number.
Return the best-guess formula that gives the force F as a function of the inputs.
Example signature: def discovered_law(mass1, mass2, distance):
  DataBlind Response by GPT-4.1 (Wrong)
def discovered_law(mass1, mass2, distance):
    G = 1  # altered universe gravitational constant
    return G * (mass1 * mass2) / (distance**3)
  Ground-truth Target Law
HIDDEN_CONSTANT_C * (mass1 * mass2) ** 2 / distance ** 1.5
@@ E.4 Exploration and Exploitation Tokens
E.5 Symbolic Regression Baselines
For comprehensive analysis, we conducted further experiments on existing methods in symbolic regression: PySR and LLM-SR, within our benchmark. As law discovery from systems (where target law is embedded in model systems) is not supported by symbolic regression methods, we only tested it in our vanilla equation setting. Below, we report and analyze the results.
PySR
PySR (cranmer2023interpretablemachinelearningscience) is an open-source Python library for symbolic regression that utilizes evolutionary algorithms to discover interpretable mathematical equations from data. We evaluated PySR across all 12 domains and 3 equation difficulty levels under the vanilla equation setting, with results illustrated in Figure 21. We observe that, in terms of symbolic accuracy, PySR yields performance comparable to or better than non-reasoning LLMs, while remaining significantly below all reasoning models. However, benefiting from its ability to conduct iterated regression, PySR achieved stronger relative performance, comparable to reasoning models in terms of RMSLE. Moreover, on easy equations, PySR outperformed all LLMs except Gemini-2.5-Pro in RMSLE, but it drops to fifth-best when facing hard equations. This indicates that this SR approach is fragile when discovering complex equations.
Figure 21: Performance of PySR in NewtonBench (vanilla equation).
LLM-SR
LLM-SR (shojaee2025llmsrscientificequationdiscovery) is a symbolic regression framework that utilizes LLM as the foundation for equation discovery, supported by automated regression tools. In this study, we examine its performance after 2,500 iterations as recommended in the original paper [^4] As LLM-SR is extremely resource-intensive (one LLM-SR inference requires 10,000 API calls, 1,000 times more than NewtonBench’s Agent), we limit our testing to the simplest domain (Law of Sound Speed)..
Table 21 presents the experimental results of LLM-SR utilizing GPT-4.1-mini as the backbone. LLM-SR marginally surpasses PySR in both accuracy and RMSLE, a finding consistent with the results reported in the original study. However, despite requiring significantly higher computational resources, LLM-SR still underperforms our code-assisted agent in terms of accuracy.
Table 21: Performance Comparison of LLM-SR. *SA criteria for LLM-SR were relaxed. As raw outputs exhibited numerical and structural noise resulting in a zero score by the default judge, post-processing was applied (rounding to four decimal places and structural simplification).
Method LLM Backbone Symbolic Accuracy (SA) % RMSLE
easy medium hard Avg.
PySR 83.3 25.0 0.0 36.1 $2.51\times 10^{-4}$
LLM-SR* GPT-4.1-mini 80.0 66.7 20.0 55.6 $\mathbf{3.89\times 10^{-5}}$
Vanilla Agent (newtonbench) GPT-4.1-mini 41.7 0.0 0.0 13.9 2.27
Agent w/ Code Asst. (newtonbench) GPT-4.1-mini 100.0 100.0 25.0 75.0 $4.61\times 10^{-2}$
Generally, symbolic regression methods leverage iterative regression and optimization to achieve superior RMSLE scores. Nevertheless, they remain sensitive to equation complexity and underperform our agent in symbolic accuracy, even while utilizing substantially greater computational resources.
E.6 Further Discussions
In this section, we further discuss NewtonBench’s implication in general scientific discovery and future directions of benchmarking.
Implications for Scientific Discovery.
NewtonBench evaluates the core cognitive architecture required for empirical research, mirroring the interplay of inductive synthesis and deductive verification in classical physics. The benchmark captures the intelligence displayed in foundational breakthroughs, such as J.J. Thomson’s discovery of the electron via cathode ray manipulation. In this classical paradigm, discovery relies on the active isolation of variables to distill symbolic laws from physical behavior. This stands in contrast to the data-driven paradigm of modern High-Energy Physics (HEP)—exemplified by the statistical reconstruction of the Higgs boson from petabytes of collision data at the Large Hadron Collider (LHC)—which often necessitates high-dimensional analysis and probabilistic inference. Nevertheless, a unified epistemic challenge holds across both eras: the necessity to abstract parsimonious mathematical descriptions from observation. By simulating this fundamental loop, NewtonBench positions itself not merely as a puzzle solver, but as a rigorous test of the reasoning engine required to recover the “source code” of physical phenomena.
However, it is important to note that while NewtonBench provides the experimental environment, real-world research necessitates the design of the experiment itself. In authentic scientific inquiry, the interaction space is rarely pre-defined; researchers must conceive and construct the apparatus required to isolate a specific phenomenon. Consequently, the current evaluation focuses on the strategic manipulation of variables and the inference of laws within a bounded system, rather than the creative genesis of the experimental setup. While this limits the scope regarding the full pipeline of autonomous discovery, it allows for a precise measurement of the agent’s ability to bridge the gap between tabletop experimentation and theoretical abstraction.
Future Directions in Discovery Benchmarking.
To bridge the gap between this foundational reasoning test and the complexity of real-world scientific discovery, future benchmarks must embrace greater open-endedness. We envision the next generation of evaluations evolving along three critical dimensions:
  • First-Principles Model Construction: Agents must advance beyond merely fitting parameters within a predefined symbolic grammar. Instead, they should be challenged to construct the model structure itself, deriving governing differential equations or symbolic laws directly from first principles.
  • Stochastic & Constrained Experimentation: Benchmarks must simulate the practical “friction” of physical science. This requires agents to maximize information gain under strict budgets while navigating imperfect data environments characterized by instrument noise, stochasticity, and confounding factors.
  • Operational Laboratory Execution: Data acquisition should evolve from simple function calls to complex procedural execution. Agents must reason about and manipulate the apparatus itself—ranging from dry lab tasks like debugging simulation pipelines to wet lab challenges like controlling robotic hardware for optical alignment—to create the conditions necessary for valid observation.
By establishing NewtonBench as a controlled starting point, we aim to guide the community toward developing agents capable of navigating these increasingly demanding frontiers.
E.7 Use of Large Language Models